Cullumber Company has provided information on intangible assets as follows.

A patent was purchased from Marin Company for $1,800,000 on January 1, 2019. Cullumber estimated the remaining useful life of the patent to be 10 years. The patent was carried in Marin’s accounting records at a net book value of $1,800,000 when Marin sold it to Cullumber.

During 2020, a franchise was purchased from Bramble Company for $510,000. In addition, 5% of revenue from the franchise must be paid to Bramble. Revenue from the franchise for 2020 was $2,530,000. Cullumber estimates the useful life of the franchise to be 10 years and takes a full year’s amortization in the year of purchase.

Cullumber incurred research and development costs in 2020 as follows.
Materials and equipment

$159,000

Personnel

195,000

Indirect costs

106,000

$460,000


Cullumber estimates that these costs will be recouped by December 31, 2023. The materials and equipment purchased have no alternative uses.

On January 1, 2020, because of recent events in the field, Cullumber estimates that the remaining life of the patent purchased on January 1, 2019, is only 5 years from January 1, 2020.

Collapse question part

(a)

Prepare the intangibles section of Cullumber’s balance sheet at December 31, 2020. (Enter account name only and do not provide descriptive information.)
CULLUMBER COMPANY
Intangibles Section of Balance Sheet

choose the accounting period

For the Month Ended December 31, 2020December 31, 2020For the Year Ended December 31, 2020

select an opening section name

Current AssetsCurrent LiabilitiesIntangible AssetsLong-term InvestmentsLong-term LiabilitiesProperty, Plant and EquipmentStockholders' EquityTotal AssetsTotal Current AssetsTotal Current LiabilitiesTotal Intangible AssetsTotal LiabilitiesTotal Liabilities and Stockholders' EquityTotal Long-term InvestmentsTotal Long-term LiabilitiesTotal Property, Plant and EquipmentTotal Stockholders' Equity

enter a balance sheet item

$enter a dollar amount

enter a balance sheet item

enter a dollar amount

select a closing section name

Current AssetsCurrent LiabilitiesIntangible AssetsLong-term InvestmentsLong-term LiabilitiesProperty, Plant and EquipmentStockholders' EquityTotal AssetsTotal Current AssetsTotal Current LiabilitiesTotal Intangible AssetsTotal LiabilitiesTotal Liabilities and Stockholders' EquityTotal Long-term InvestmentsTotal Long-term LiabilitiesTotal Property, Plant and EquipmentTotal Stockholders' Equity

$enter a total amount for this section
Answer #1
Intangible Assets
Patents from Marin company 1,296,000
Franchise from Bramble Company 459,000
Total intangible assets 1,755,000

Amortization of patent for 2019 = 1,800,000/10 = 180,000

Book value of patent at the end of 2019 = (1,800,000-180,000) = 1,620,000

Amortization of patent for 2020 = (1,620,000/5) = 324,000

Book value of patent at the end of 2020 = 1,620,000 - 324,000 = 1,296,000